OpenStudy (anonymous):

How do I figure out if these are Rational and Irrational? 1. -101 2. 21.192 3. Square Root of 9 4. 8/17 5. Pie 6. Square Root of 2 7. 39.81 (there's a line over 81) @Hero @mathslover @mathlover2014 @mathhelpwanted @mathstudent55 @jim_thompson5910 I need you to explain it really simple, because I really don't understand it. Well...I understand it to an extent. Don't worry this isn't a test or homework. Someone gave me some stuff to practice and I want to learn it to be ready for high school. :)

OpenStudy (anonymous):

I'll be right back, I have to eat dinner.

hero (hero):

A number is rational if it can be written in the form a/b, where a and b are integers. Otherwise it is irrational.

OpenStudy (anonymous):

Like @Hero said, a number is rational if it can be written in the form a/b. In other words, if it can be written as a fraction. Numbers like pi are irrational because they are infinite and cannot be put into fractional form. It can be approximated to 22/7, but that is not exact.

OpenStudy (anonymous):

OK, Can you walk me through each one?

OpenStudy (anonymous):

@girlnotonfire Can you walk me through each one? :)

OpenStudy (anonymous):

-101 would be.... -101/1?

OpenStudy (anonymous):

Yes, so it is rational.

OpenStudy (anonymous):

21.192 since the decimal expansion isn't made up of zeros or a single digit repeating is and irrational number?

OpenStudy (anonymous):

21.192=21+192/1000=21+24/125=2649/125

hero (hero):

21.192 = 21 + 192/1000 = 21000/1000 + 192/1000 = 21192/1000 Therefore 21.192 is rational

OpenStudy (anonymous):

OK, so that means it's Rational?

OpenStudy (anonymous):

Yes, it is rational.

OpenStudy (anonymous):

Oh, I get it thanks! Hold On and let me see if I can figure out the answer for the next one. :)

OpenStudy (anonymous):

The Square root of 9 is rational?

OpenStudy (anonymous):

Yes! The square root of 9 is 3 or 3/1. Good job :)

OpenStudy (anonymous):

Now 8/17 is already a fraction, so how do I work that? @girlnotonfire

OpenStudy (anonymous):

If it is already a fraction, we automatically know it is rational, no work needed :)

OpenStudy (anonymous):

OK, :) Now the next one look like pie, or n but it's completely straight on both side and straight on top. I don't know what that is. Do you? 0_0 @girlnotonfire

OpenStudy (anonymous):

That's probably pi as well, I'm not sure what else it could be. If it looks like Π or π, it is definitely pi.

OpenStudy (anonymous):

It looks like the first one, which if you said it was pie.. Then it's irrational. Right? :) @girlnotonfire

OpenStudy (anonymous):

Yes, because it is an infinite decimal that cannot be made into an exact fraction.

OpenStudy (anonymous):

I already know that #6 is irrational. :) So, that leaves #7. Since there is a line over the 81 in 39.81 it means it repeats like this.... 39.81818181 does that mean it's Rational?

OpenStudy (anonymous):

Yes, all repeating decimals are classified as rational numbers. In this case, 39.81(repeating)=438/11, so it is rational.

OpenStudy (anonymous):

Thanks! :)